You are given the following information concerning a stock and the market:

Returns
Year Market Stock
2011 20 % 32 %
2012 10 10
2013 20 4
2014 −15 −29
2015 37 16
2016 15 16

a. Calculate the average return and standard deviation for the market and the stock. (Use Excel to complete the problem. Do not round intermediate calculations. Enter your answers as a percent rounded to 2 decimal places.)

b. Calculate the correlation between the stock and the market, as well as the stock’s beta. (Use Excel to complete the problem. Do not round intermediate calculations. Round your correlation answer to 2 decimal places and beta answer to 4 decimal places.)
Answer #1

* x NI : xv fx A B C D E F G H I J K L M N O P A 1 2 3 a) Market Stock b c = b - 14.5 e f g=f-8.17 i=ctg Return(M)- Return(M)

... X N45 x V fic A B C D E F G H I J K L M N O P Standard Deviation of Market = (1457.5 / 6)^(1/2) SD(M) = 15.59% Standard D